Curriculum Vitae

Dr. Marcelo de Lima Leal Ferreira

Present position

Period 2014 – Present
Institute Leiden Observatory  The Netherlands
Position Research Associate
Fellowship CNPq Science without Borders/Brazil (2014–present)
Group Sterrewacth Leiden ISM Group (head: Prof. Xander Tielens)

Education

Period 2010 – 2014
Institute Argelander-Institut für Astronomie/UniBonn Germany
Degree PhD
Fellowship DAAD/Germany (2011–2014); contract (2010–2011)

Project Magnetic Fields and the Formation of Aspherical of Planetary Nebulae (Advisor: Dr. Wouter Vlemmings)

Period 2007 – 2009
Institute Observatório do Valongo/UFRJ  Brazil
Degree Master
Fellowship CAPES/Brazil

Project Physico-chemical Spectroscopic Mapping of the Planetary Nebula NGC 40 (Advisors: Dr. Denise Gonçalves & Dr. Hektor Monteiro)

Period 2000 — 2007
Institute Observatório do Valongo/UFRJ  Brazil
Degree Bachelor
Fellowship CNPq/Brazil & FAPERJ/Brazil

Project Wavelet Technique Applications on Planetary Nebulae Images (Advisor: Dr. Carlos Rabaça)

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2

Complementary Education on Astrophysics

Period 2011 – 2011
Institute INAF – Institute of Radioastronomy (Italy)
Course European Radio Interferometry School (ERIS)
Period 2009 – 2009
Institute Instituto Nacional de Pesquisas Espaciais (Brazil)

Course III INPE ADVANCED COURSE ON ASTROPHYSICS: ASTROSTATISTIC

Period 2008 – 2008
Institute Observatório Nacional (Brazil)
Course XIII Cicle of Special Courses
Period 2007 – 2007
Institute Universidad de Concepción (Chile)
Course Third Chilean Advanced School on Astrophysics
Period 2007 – 2007
Institute Observatório Nacional (Brazil)
Course XII Cicle of Special Courses
Period 2006 – 2006
Institute Observatório Nacional (Brazil)
Course XI Cicle of Special Courses
Period 2005 – 2005
Institute European Southern Observatory (Chile)
Course Latin American Astronomy Summer School
Period 2004 – 2004
Institute Observatório Nacional (Brazil)
Course IX Cicle of Special Courses
Period 2004 – 2004
Institute Centro Brasileiro de Pesquisas Físicas (Brazil)
Course Nuclear Physics and Stellar Evolution − V School
Period 2004 – 2004
Institute Centro Brasileiro de Pesquisas Físicas (Brazil)
Course Image Processing − V School

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2

Period 2004 – 2004
Institute Observatório do Valongo / UFRJ (Brazil)
Course Stellar Formation on Galaxies − I GEMAC Course
Period 2003 – 2003
Institute Observatório Nacional (Brazil)
Course VIII Cicle of Special Courses
Period 2000 – 2000
Institute Fundação Planetário da Cidade do Rio de Janeiro (Brazil)
Course Sky Identification

Languages

Portuguese Good conversation, reading and writing (First language)
English Good conversation, reading and writing
Spanish Moderate conversation; basic reading
German Basic conversation; basic reading

Awards, Fellowships & Grants

2016 Individual 50% grant for IAUS323 Conference (IAU)
2016 Individual 50% grant for IAUS323 Conference (LKBF)
2015 Fellowship Science without Borders (1 year extension) (CNPq)
2014 Fellowship Science without Borders (1 year) (CNPq)
2014 PhD Fellowship (6 months extension) (DAAD)
2013 PhD Fellowship (1 year extension) (DAAD)
2012 PhD Fellowship (1 year extension) (DAAD)
2011 PhD Fellowship (1 year) (DAAD)
2008 Academic Highlight Award (CCMN/UFRJ)
2007 Masters Fellowship (2 years) (CNPq)

2007 Bachelor Research Fellowship (extension until interruption at graduation) (CNPq/PIBIC)

2007 Individual 50% grant for APN4 Conference (APN4 SOC)
2007 Individual 50% grant for LARIN2007 Conference (IAU)
2006 Bachelor Research Fellowship (1 year) (CNPq/PIBIC)
2006 Best talk in Sputnik Session (Scientific Initiation Journey, UFRJ)
2004 Bachelor Research Fellowship (1 year extension) (FAPERJ)
2003 Bachelor Research Fellowship (1 year) (FAPERJ)

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2 2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s m l l f e r r e i r a @ g m a i l . c o m

+ 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2

Scientific Production

LIST OF PUBLICATIONS – REFEREED

2016 “XSHOOTER spectroscopy of the enigmatic planetary nebula Lin49 in the Small Magellanic Cloud” (Otsuka, M.; Kemper, F.; Leal-Ferreira, M. L.; Aleman, I.; Bernard-Salas, J.; Cami, J.; Ochsendorf, B.; Peeters, E.; Scicluna, P., MNRAS, 462, 12)

2015 “Magnetic Fields and the Formation of Aspherical Planetary Nebulae” (Leal-Ferreira, M. L., PhD Thesis, Universität Bonn, urn:nbn:de:hbz:5n-39033)

2014 “The detached dust shells around the carbon AGB stars R Sculptoris and V644 Scorpii” (Maercker, M.; Ramstedt, S.; Leal-Ferreira, M. L.; Olofsson, G.; Floren, H. G., A&A, 570, 101)

2014 “The wonderful complexity of the Mira AB system” (Ramstedt, S.; Mohamed, S.; Vlemmings, W. H. T.; Maercker, M.; Montez, R.; Baudry, A.; De Beck, E.; Linqvist, M.; Olofsson, H.; Humphreys, E. M. L.; Jorissen, A.; Kerschbaum, F.; Mayer, A.; Wittkowski, M.; Cox, N. L. J.; Lagadec, E.; Leal-Ferreira, M. L.; Paladini, C.; Perez-Sanchez, A.; Sacuto, S., A&A, 570, 14)

2013 “Spatially resolved physical and chemical properties of the planetary nebula NGC 3242” (Monteiro, H.; Gonçalves D. R.; Leal-Ferreira, M. L.; Corradi, R. L. M., A&A, 560, 102)

2013 “Magnetic fields around evolved stars: further observations of H2O maser polarization” (Leal-Ferreira, M. L.; Vlemmings, W. H. T.; Kemball, A.; Amiri, N., A&A, 554, 134)

2012 “Rotten Egg nebula: the magnetic field of a binary evolved star” (Leal-Ferreira M. L., Vlemmings W. H. T., Diamond P. J., Kemball A., Amiri N., Desmurs, J.-F., A&A, 540, 42)

2011 “Physico-chemical spectroscopic mapping of the planetary nebula NGC 40 and the 2D NEB, a new 2D algorithm to study ionized nebulae” (Leal-Ferreira M. L., Gonçalves D. R., Monteiro H., Richards J. W., MNRAS, 411, 1395)

2009 “Caracterização físico-química da nebulosa planetária NGC 40 através de mapeamento espectroscópico” (Leal-Ferreira, M. L., Master Disseration, Universidade Federal do Rio de Janeiro

2007 “Aplicações da Tćnica de Wavelet a Imagens de Nebulosas Planetárias” (Leal-Ferreira, M. L., Bachelor Project, Universidade Federal do Rio de Janeiro

LIST OF PUBLICATIONS – MAGAZINES

2015 Maercker, M.; Ramstedt, S.; Leal-Ferreira, M. L.; Olofsson, G.; Floren, H. G; The ESO Messenger, 160, 9

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2

LIST OF PUBLICATIONS – CONFERENCE PROCEEDINGS

2016 Otsuka, M; Kemper, F.; Leal-Ferreira M.; Aleman, I.; Bernard-Salas, J.; Cami, J.; Ochsendorf, B.; Peeters, E., JPhCS, 728e2006O

2014 Leal-Ferreira M.; Vlemmings, W. H. T.; Kemball, A.; Amiri, N.; Maercker, M.; Ramstedt, S.; Olofsson, G., apn6, 49

2012 Monteiro H., Gonçalves D., Leal-Ferreira M., Corradi R., Sánchez S., IAUS, 283, 448

2012 Leal-Ferreira M. L., Vlemmings W. H. T., Diamond P. J., Kemball A, Amiri N., Desmurs J.-F., IAUS, 283, 418

2012 Leal-Ferreira M. L., Vlemmings W. H. T., Diamond P. J., Kemball A., Amiri N., Desmurs J.-F., IAUS, 287, 79

2011 Leal-Ferreira M. L., Gonçalves D., Monteiro H., apn5, 47
2011 Gonçalves D., Leal-Ferreira M. L., Monteiro H., apn5, 139

2009 Leal Ferreira M. L., Rabaca C. R., Cuisinier F., Epitácio Pereira D. N., RMxAC, 35, 290 (Larin 2007)

2009 Leal Ferreira M. L., Rabaca C. R., Cuisinier F., Moisés A. P., Epitácio Pereira D. N., Asymmetrical Planetary Nebula IV, p. 211−216

2009 Leal Ferreira M. L., Rabaca C. R., Cuisinier F., Epitácio Pereira D. N., Asymmetrical Planetary Nebula IV, p. 217−222

2005 Cuisinier F., Moisés A. P., Fereira M. L., Rabaca C. R., Gonçalves D. R., Planetary Nebulae as Astronomical Tools, p. 28−31

Principal Investigator Approved Observational Proposals

Title The Magnetic Field Morphology Around Evolved Stars
Telescope Awarded with 3 runs at eMERLIN

Student Supervisions

2016 Co-advisor of Master Student Esmee Stoop, “Double Bubble Toil and Trouble”, Leiden Obsertory, The Netherlands

2016 Co-advisor of Master Student Andrew Barr, “Analysis of the dust content of (ultra) compact HII regions from FORCAST/SOFIA images. Comparison of the results with those of Orion and W3A”, Leiden Obsertory, The Netherlands

2016 Co-advisor of Summer Student Soheb Mandhai, “Cannibalism in stars: progenitors of supernovae”, Leiden Obsertory, The Netherlands

2015 Main advisor of Summer Student Andrea Gaughan, “The Fellowship of the Ring Nebula: The North Tower of Gemini”, Leiden Obsertory, The Netherlands

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2 2015 Main advisor of Summer Student Samuel Kerber, “The Fellowship of the Ring Nebula: The Return of the King HST”, Leiden Obsertory, The Netherlands

Symposiums, Conferences, Meetings, and Events

2016 IAUS 323 - Planetary nebulae: Multi-wavelength probes of stellar and galactic evolution (Beijing, China) – Contribution: Poster

2016 NOVA PhD/Postdoc Weekend (Amersfoort, The Netherlands)

2016 WORKPLANS - WORKshop for PLAnetary Nebulae observationS (Leiden, The Netherlands) – Main-organizer

2015 30 Years of Photoionization Regions (Asilomar, USA) – Co-organizer

2015 225th Meeting of the American Astronomical Society (Seattle, USA)

2014 DAAD Orientierungsseminar (Bonn, Germany) – Contribution: Invited Talk

2013 Asymmetrical Planetary Nebulae VI (Riviera Maya, Mexico) – Contribution: Talk

2013 DAAD Orientierungsseminar (Oct/2013) (Bonn, Germany) – Contribution: Invited Talk

2013 DAAD Orientierungsseminar (May/2013) (Bonn, Germany) – Contribution: Invited Talk

2012 Young European Radio Astronomy Conference (Pushchino, Russia)

– Contribution: Talk

2012 Evolved stars with ALMA Cycle 1 (Onsala, Sweden)

2012 DAAD Orientierungsseminar (Bonn, Germany) – Contribution: Invited Talk

2012 IAU Symposium 287: Cosmic masers - from OH to H0 (Stellenbosch, South Africa) – Contribution: Poster

2011 IAU Symposium 283: Planetary Nebulae - an Eye to the Future (Tenerife, Spain) – Contribution: Poster

2011 Evolved Stars with ALMA early Science (Bonn, Germany)
2011 German ALMA Early Science Community Day (Bonn, Germany)
2011 DAAD Stipendiatentreffen (Dortmund, Germany) – Contribution: Talk
2010 Training for the usage of the AstroBEAR code (Rochester, USA)

2010 Why Galaxies Care About AGB Stars II - Shining Examples and Common Inhabitants (Vienna, Austria) – Contribution: Poster

2010 Asymmetric Planetary Nebulae V (Windermere, England) – Contribution: Poster

2009 IAU XXVII General Assembly (Rio de Janeiro, Brazil) – Contribution: Poster

2009 Week of the Grad Students of the Observatório do Valongo (Rio de Janeiro, Brazil) – Contribution: Talk

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2 2008 50 Years of the Astronomy Undergraduation Course (Rio de Janeiro, Brazil) – Contribution: Talk

2007 Asymmetrical Planetary Nebulae IV (La Palma, Spain) – Contribution: Poster

2007 Annual Meeting from the Brazilian Astronomical Society (Passa Quatro, Brazil) – Contribution: Talk

2007 XII Latin-American Regional Meeting of the IAU (Isla Margarita, Venezuela) – Contribution: Poster

2007 XXIX Scientific Initiation Journey (Rio de Janeiro, Brazil) – Contribution: Talk

2006 XXXII Annual Meeting from the Brazilian Astronomical Society (Atibaia, Brazil) – Contribution: Poster

2006 First Brazilian Workshop on Astrobiology (Rio de Janeiro, Brazil))

2006 XXVIII Scientific Initiation Journey (Rio de Janeiro, Brazil) – Contribution: Talk

2006 National Week of Science and Technology (Outreach) (Rio de Janeiro, Brazil)

2005 XXXI Annual Meeting from the Brazilian Astronomical Society (Águas de Lindóia, Brazil) – Contribution: Poster

2005 XXVII Scientific Initiation Journey (Rio de Janeiro, Brazil) – Contribution: Talk

2005 National Week of Science and Technology (Outreach) (Rio de Janeiro, Brazil)

2005 Astronomy at Casimiro de Abreu (Outreach) (Casimiro de Abreu, Brazil)

2005 UFRJmar (Outreach) (Arraial do Cabo, Brazil)

2004 XXX Annual Meeting from the Brazilian Astronomical Society (São Pedro, Brazil) – Contribution: Poster

2004 XXVI Scientific Iniciation Journey (Rio de Janeiro, Brazil) – Contribution: Talk

2004 National Week of Science and Technology (Outreach) (Rio de Janeiro, Brazil)

2004 Integrated Knowledge Workshop (Outreach) (Arraial do Cabo, Brazil)

2003 XXIX Annual Meeting from the Brazilian Astronomical Society (São Pedro, Brazil) – Contribution: Poster

2003 XXV Scientific Iniciation Journey (Rio de Janeiro, Brazil) – Contribution: Talk

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2

2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s

m l l f e r r e i r a @ g m a i l . c o m + 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2

Organization of Events

2016 WORKPLANS - WORKshop for PLAnetary Nebulae observationS (Leiden, The Netherlands) – Main-organizer

2015 30 Years of Photoionization Regions (Asilomar, USA) – Co-organizer

Invited Institute Talks

2016 Magnetic Fields and the Formation of Aspherical Planetary Nebulae (Observatório do Valongo, Rio de Janeiro, Brazil)

2011 The Magnetic Field of the Rotten Egg Nebula (Observatório do Valongo, Rio de Janeiro, Brazil)

Outreach Activities

2006 Semana Nacional de Ciência & Tecnologia, Rio de Janeiro, Brazil
2005 Astronomia em Casimiro de Abreu, Casimiro de Abreu, Brazil
2005 Semana Nacional de Ciência & Tecnologia, Rio de Janeiro, Brazil
2005 UFRJ mar, Cabo Frio, Brazil
2004 Feira do Conhecimento Integrado, Arraial do Cabo, Brazil
2004 Mostra de Extensˆ ao: Projeto da UFRJ nas Escolas, Búzios, Brazil
2004 Semana Nacional de Ciência & Tecnologia, Rio de Janeiro, Brazil

L e i d e n O b s e r v a t o r y – N i e l s B o h r w e g 2 2 3 3 3 C A L e i d e n , T h e N e t h e r l a n d s m l l f e r r e i r a @ g m a i l . c o m

+ 3 1 ( 0 ) 6 1 4 2 4 3 1 1 2